OpenStudy (anonymous):

I really need help! Differentiate: y=sin(2x+sinx) & y=tan(e^x)

OpenStudy (anonymous):

y=-cos(2x+sinx)+sin(x_cosx) y=tan(e^x)=(-cos(e^x)*e^x * cos(e^x) - sinx(e^x)*(sin(e^x)*e^x) )/(cos(e^x)*cos(e^x))